# The lonely runner conjecture holds for eight runners

### v1

@@ 1 Introduction
The Lonely Runner Conjecture is a well-known open problem in combinatorial number theory and Diophantine approximation. It was introduced independently by Wills [21] and Cusick [6], and has since attracted significant attention. Consider $k+1$ runners with distinct constant speeds running around a unit-length circular track. The lonely runner conjecture asserts that for any runner there is a time at which the runner is at least a distance $1/(k+1)$ away from every other runner. The conjecture can be stated more formally as follows.
@@
Other formulations of the problem allow $v_{1},\ldots,v_{k}$ to be real numbers, but Wills proved that the two conjectures are equivalent, and we can focus our attention on the integral case [22]. The name and runner interpretation of this conjecture is due to Goddyn [4]. Adding the same constant to all speeds does not change the instance of the problem, so one may, without loss of generality, set the speed of one runner to 0. Moreover, if one only cares about the distance to 0, then the speeds $x$ and x-x play symmetric roles, which implies the following equivalent formulation of the conjecture.
@@
Some partial results on the Lonely runner conjecture in the context of lacunary sequences were also obtained by different authors [2, 10, 11, 14, 19]. A stronger version of these results can easily be deduced as a corollary of a recent result on lacunary sequences [18]. In Section 8; we present the context, and we provide the improved result.

### v2

@@ 1 Introduction
The Lonely Runner Conjecture is a well-known open problem in combinatorial number theory and Diophantine approximation. It was first introduced by Wills [21, 22] with a purely number theoretic formulation and has since attracted significant attention. Cusick proposed another formulation of the problem in terms of view obstructions [6]. The following runner interpretation and current name of the conjecture is due to Goddyn [4]. Consider $k+1$ runners with distinct constant speeds running around a unit-length circular track. The lonely runner conjecture asserts that for any runner there is a time at which the runner is at least a distance $1/(k+1)$ away from every other runner. The conjecture can be stated more formally as follows. For any real $x$, $\|x\|$ is the distance from $x$ to the nearest integer.
@@
Other formulations of the problem allow $v_{1},\ldots,v_{k}$ to be real numbers, but Wills proved that the two conjectures are equivalent, and we can focus our attention on the integral case [23]. Moreover, adding the same constant to all speeds does not change the instance of the problem, so one may, without loss of generality, set the speed of one runner to 0. Finally, if one only cares about the distance to 0, then the speeds $x$ and x-x play symmetric roles, which implies the following equivalent formulation of the conjecture.
@@
Some partial results on the Lonely runner conjecture in the context of lacunary sequences were also obtained by different authors [2, 10, 11, 14, 19]. A stronger version of these results can easily be deduced as a corollary of a recent result on lacunary sequences [18]. In Section 8, we present the context, and we provide the improved result.